The Griggs Corporation has credit sales of $1,075,900 Total assets turnover Cash to total assets Accounts receivable turnover Inventory turnover Current ratio Debt to total assets 2.90 times 1.30 % 10 times 14 times 1.88 times 45 % Using the above ratios, complete the balance sheet. (Round your answers to the nearest whole dollar.) GRIGGS CORPORATION Balance Sheet Assets Liabilities and Stockholders Equity Cash Accounts receivable Inventory Current debt Long-term debt Total debt Total current assets Fixed assets Equity Total assets lotal debt and stockholders equity
Answer #1
GRIGGS CORPORATION
Balance Sheet
Assets Liabilities and stockholders equity
Cash $4,823.00 Current Debt $100,672.00
Accounts Receivables $107,590.00 Long term debt $66,278.00
Inventory $76,850.00 Total Debt $166,950.00
Total Current Assets $189,263.00
Fixed Assets $181,737.00 Equity $204,050.00
Total Assets $371,000.00 Total Debt and Stockholders Equity $371,000.00
Working
Total assets turnover = Sales / Total Assets = 2.90 times
Hence , Total Assets = Sales / 2.90 = $1075900 / 2.90 = $3,71,000
Cash to total assets = Cash / Total Assets = 1.30%
Hence , Cash = Total Assets * 1.30% = $371000 * 1.30% = $4823
Accounts receivable turnover = Sales / Accounts Receivables = 10 times
Hence , Accounts Receivables = Sales / 10 = $1075900 / 10 = $1,07,590
Inventory turnover = Sales / Inventory = 14 times
Hence , Inventory = Sales / 14 = $1075900 / 14 = $76,850
Total Current assets = Cash + Accounts receivables + Inventory = $189263
Fixed Assets = Total assets - Total current assets = $371000 - $189263 = $1,81,737
Current ratio = Current assets / Current Liabilities = 1.88 times
Hence , Current liabilities = Current assets / 1.88 = $189263 /1.88 = $1,00,672
Debt to total assets = Total Debt / Total assets = 45%
Total debt = Total assets * 45% = $371000 * 45% = $1,66,950
Long term debt = Total debt - Current debt = $66,278
Equity = Total equity and debt - Total debt = $371000 - $166950 = $204050